Equal weights of ethane and hydrogen are mixed in an empty container at 25°C. the fraction of the total pressure exerted by hydrogen is
a. 1 : 2
b. 1 : 1
c. 1 : 16
d. 15 : 16

Grade:11

3 Answers

Kevin Nash
9 years ago
Sol. Let x g of each gas is mixed. Mole of ethane = x/30 Mole of hydrogen = x/2 Mole fraction of hydrogen = (x/2)/(x/2 + x/30) = 15/16 ⇒ (Partical pressure of H_2)/(Total pressure) = mole fraction of hydrogen = 15 : 16
Nutan Raj
4 years ago
given,equal wt of ethane& hydrogen are mixed in an empty container at 25 degree celcius.
Initial gram wt of C2H6=wg
Initial gram wt of H2=wg
No of moles of C2H6=w÷30
No of moles of H2=w÷2
According to Henry law
PH2/Ptotal=XH2=nH2/nH2+nC2H6=w/2÷w/2+w/30
PH2/Ptotal=15/16
ankit singh
3 years ago

Let x grams of each gas is mixed.


A Mole of ethane is x/30.
A mole of hydrogen is x/2.
Mole fraction of hydrogen is 2x+30x2x=1615
 
Partial pressure of hydrogen/total pressure = mole fraction of hydrogen =15:16